# Localized Multi-Dimensional Patterns

### v1

@@ 2.4.4 Oscillons and Breathers
While all the structures we have looked at so far are stationary, time-periodic spatially localized structures have been extensively studied in a wide-range of models. These structures are called either oscillons or breathers, depending on the context they arise in. Oscillons can arise in nonlinear systems from a uniform background state under parametric forcing whose frequency is often close to the frequency of the pattern oscillation in space (the wavenumber of the rolls) [5, 6, 164]. Such localized time-periodic patterns have been observed in Newtonian fluids [11, 101, 242, 262], chemical reactions [195, 238, 236], granular layers [231], and colloidal suspension [152].
@@
In the framework of (2.27), oscillons correspond to localized solutions that are time-independent. The work [57] demonstrated that there are two different types of localized steady-states in (2.27), as depicted in Figure 14. The first is the usual localized solution that limits to $U=0$ as $x\to\pm\infty$. The second limits to a nontrivial background state $U_{\mathrm{unif}}\neq 0$ as $x\to\pm\infty$ and were referred to as reciprocal oscillons, also superoscillons in earlier work [33]. Both of these types of oscillons can have either decay monotonically to $U=0$ as $x\to\pm\infty$, leading to spatial profiles resembling Gaussians/solitons, or damped oscillations into $U=0$, leading to spatial profiles resembling the localized rolls of the SHE.
@@ 2.4.5 Transition to the Singular Limit
Figure 16: Spatial eigenvalues of the quiescent state for the Belyakov–Devaney transition
@@ 2.4.6 Invasion Fronts Outside the Snaking Region
Figure 17: Outside of the snaking parameter regime localized roll patterns are dynamic, growing for $\mu$ to the left (red) of the snaking branches and collapsing in on itself to the right (blue). This growth is non-uniform, alternating between fast and slow expansions, depending on if the pattern is close to a steady-state at a nearby saddle-node.
Throughout this review we primarily focus on steady-state patterns. However, some work has been done to understand dynamic localized patterns. For example, outside of the snaking parameter regime in the SHE one can find that either the rolls expand into the background state or collapse upon themselves as time progresses, as illustrated in Figure 17. Despite the snaking region of parameter space typically being much smaller than the invasion regions, significantly less is known about these patterns and processes.
The spreading of rolls into the trivial state is fundamentally different in nature to the spreading of the trivial state. That is, in the case where the trivial state invades the roll patterns, the speed of the invasion is dictated by the wavelength of the rolls themselves. However, for the roll invasion process, since we are in the bistable region, the wavelength of the deposited pattern and the time between stripes being deposited are selected [156, 9]. The roll invasion fronts do not invade at a constant speed and instead have a “helical” invasion process, coming as a more complicated version of the “bottlenecks” one encounters in parameter regimes near a saddle-node bifurcation [216, Section 4.3]. Precisely, the non-constant traveling speed of the front is understood by looking at the fronts near the snaking region, where one would expect the invasion process to slow down as one passes near a fold bifurcation and then speed up between the folds, as illustrated in Figure 17. There has been a lot of recent progress on the analysis of pattern-forming invasion fronts; see [104] for a review.
@@ 3 Localization in Only One Direction
We begin this section in §3.1 with localized stripe patterns wherein the localized pattern takes the form of a 1D roll pattern extended trivially in either the direction of $x$, $y$, or a combination thereof. §3.2 reviews the work on patterns with hexagonal patches arranged along the region of localization, while §3.3 provides the spatial dynamics interpretation of these stripe and front patterns. Finally, in §3.4 we review similar work on planar lattices.
@@ 3.3 Spatial Dynamics
Thus, we see that many of the same ingredients that were exploited in [30] (and summarized in §2.3) are present to prove the existence of localized roll patterns in 1D. Much like the 1D setting, localized structures of (3.1) correspond to orbits homoclinic to $\mathbf{0}\in\mathcal{X}$, i.e. $\mathbf{u}(x,\cdot)\to\mathbf{0}$ as $x\to\pm\infty$. One needs to be careful here though as (3.6) is ill-posed as an initial-value problem. However, the theory developed in [194, 206] guarantees that stable and unstable manifolds exist for the quiescent state $\mathbf{0}\in\mathcal{X}$, providing the necessary phase-space structure for homoclinic orbits to exist. Beyond this, first recall from §2.3 that spatial dynamics interprets localized solutions as homoclinic orbits wrapping a periodic orbit to produce the localized roll patterns. The same interpretation now holds here where the localized pattern comes from the homoclinic orbit of (3.6) wrapping around a periodic in $x$ solution $\gamma(x;y)$. Again, the theory in [206] guarantees the existence of stable and unstable manifolds to $\gamma(x;y)$ in the phase-space $\mathcal{X}$ that can be used to prove the existence of a homoclinic orbit that wraps around $\gamma$ (in $x$), producing the desired localized solution to (3.1).
@@ 4.1 Emergence of Axisymmetric Patterns
Theorem 4.1 (Spot A, [153, Theorem 2]).
@@
Theorem 4.2 (Rings, [153, Theorem 1]).
@@
Notice that these solutions bifurcate from $\mu=0$ in pairs, giving that rings arise via a pitchfork bifurcation from the Turing instability in the SHE. Moreover, the condition $|\nu|>\sqrt{27/38}$ arises from (4.6) and translates to the requirement that the rolls (Turing patterns) in the SHE bifurcate subcritically. In what follows we refer to solutions along the bifurcating $u_{R}(r)$ branch as Ring A and the solutions along the $-u_{R}(r)$ branch as Ring B.
@@
Theorem 4.3 (Spot B, [171, Theorem 1]).
@@ 4.2 Axisymmetric Snaking Branches
Continuing the axisymmetric patterns of the previous subsection from $\mu\approx 0$ into larger values of $\mu$ reveals a complex bifurcation structure that bears little resemblance to the familiar 1D snakes and ladders profiles [170]. The situation appears to be qualitatively the same for both dimensions $n=2,3$ and Figure 29 presents a numerically obtained bifurcation diagram for visual reference. Initially the Spot A and Spot B branches follow the usual snaking structure wherein fold bifurcations flip back and forth over a compact region of parameter space with the folds appearing to align along two vertical asymptotes. Moreover, like the 1D snaking phenomenon, rolls are added that extend the region of localization for the axisymmetric pattern, in turn causing the $L^{2}$-norm of the solution to increase, thus resulting in the vertical ascension in the bifurcation diagram.
@@
This vertical growth along the bifurcation diagram is not long lived. Eventually both the Spot A and Spot B branches reach a maximal finite height in the bifurcation diagram where they turn into rings and begin decreasing in norm by subtracting rolls. Following [170], Spot A connects to the Ring A branch and Spot B the Ring B branch, as defined in the previous subsection. Above both of these terminal snaking branches in the bifurcation diagram lie stacks of isolas - closed curves in the bifurcation diagram. Along each isola the solution profiles change in intricate ways and switch between both ring- and spot-like patterns. In both 2D and 3D there are only finitely many isolas with the 2D patterns extending vertically in the bifurcation diagram until they have approximately 38 rolls, while in 3D the isolas terminate with fewer rolls. Finally, above the isolas there is another snaking branch which consists of two intertwined arms that appear to ascend vertically in the bifurcation diagram without bound. This snaking branch is referred to as the upper snaking branch, while the branch containing Spot A or Spot B patterns that emerges from $(U,\mu)=(0,0)$ is termed the lower snaking branch. The upper branch differs significantly from the lower branch in that rolls are added near the core of the pattern, i.e. near $r=0$, as opposed to at the tail like in 1D. Based on numerical observations, McCalla and Sandstede further conjectured in [170] that the upper branch collapses onto a single parameter value as the number of rolls in the localized pattern goes to infinity. Bramburger et al. [43] attempted to better understand the bifurcation structure of localized axisymmetric patterns in the two- and three-dimensional SHE from an analytical perspective. As in the previous section, their work focused on the radial spatial dynamical system (4.3). When $n=1$ the nonautonomous terms disappear, reverting back to the 1-dimensional setting (2.10) where all dynamics take place inside the level sets of the Hamiltonian function (2.18). Importantly, when $n=1$ roll patterns of the SHE (4.1) correspond to periodic orbits of (4.3) which is only a slice of a ‘stack’ of periodic orbits forming a normally hyperbolic invariant manifold $\mathcal{P}$ that can be parameterized by the value of $\mathcal{H}$. The result is a cylinder of stacked periodic orbits in phase-space, as illustrated in Figure 30. Due to the conserved quantity $\mathcal{H}$, localized roll patterns of the Swift–Hohenberg equation in 1D can only be constructed through connections with the periodic orbit in $\mathcal{H}^{-1}(0)$, the zero level set of $\mathcal{H}$, to the rest state $U=0$ for each fixed $(\mu,\nu)$.
@@
Figure 30: Left: The anatomy of a radial roll solution used in the proofs of [43]. Right: In the phase-space of (4.3) localized radial roll patterns wrap around an invariant cylinder in phase-space for a finite time to produce the rolling pattern.
@@
To better understand the system (4.3) (and in turn axisymmetric localized patterns), McCalla and Sandstede proposed introducing a “dimensional perturbation” in the form of $n=1+\varepsilon$ [170]. Although meaningless in the context of the dimension of the underlying space, when $0<\varepsilon\ll 1$ this dimensional perturbation casts (4.3) as a true perturbation from the well-understood spatial system (2.10) in 1D. In particular, [43, Theorem 2.7] shows that snaking branches persist for plateau lengths $L\leq\mathrm{exp}(b/\varepsilon)$, with constant $b>0$. Such an upper bound is the result of determining the maximal number of times a trajectory can be guaranteed to wrap around the perturbed cylinder in complete generality. Surprisingly, numerical continuations of localized radial roll solutions to (4.3) at fixed values of $\varepsilon$ reveal that the upper bound $\mathrm{exp}(b/\varepsilon)$ closely tracks with the maximum of the lower snaking branches, even well beyond the asymptotic regime $0<\varepsilon\ll 1$ and into $\varepsilon=1,2$ (dimensions $n=2,3$).
@@
There are a number of questions that remain outstanding related to the bifurcation curves of localized radial roll patterns. Primary among them is that there is still no explanation for the intermediate stack of isolas that splits the lower and upper snaking branches. Similarly, it is not clear why the upper branch forms a connected curve or if it is even the uppermost branch in the bifurcation diagram. That is, it is entirely possible that the observed upper branch terminates much like the lower branch at a finite height in the bifurcation diagram, possibly giving way to more isolas or another snaking branch. Such a finding would not necessarily be surprising as the work [153] originally presented evidence that the lower branch extends indefinitely and it wasn’t until [170] that the termination of the lower branch was observed far up in the bifurcation diagram. Finally, the reason for the difficulty in predicting the specific shape of the snaking diagram with $n=2,3$ is that numerical investigations indicate that the structure is driven by the behavior of solutions away from the invariant manifold of periodic orbits. This means that the curves depend on the global dynamics of the dynamical system (4.3) and not just the local properties near the perturbed cylinder.
@@ 4.3 Time Periodic Patterns
Theorem 4.4 (Radial oscillons, [173, Theorem 1]).
@@
Figure 31: Spatial eigenvalues for the Takens–Bogdanov instability.
This result again makes use of radial spatial dynamics and provides the existence as emergence from a bifurcation. In the setting of (4.10), the curve $\gamma=\sqrt{\mu^{2}+\omega^{2}}$ corresponds to a Takens–Bogdanov bifurcation in the associated radial spatial dynamical systems; see Figure 31. The condition $\mu<\beta\omega$ provides that the bifurcation is subcritical, while we recall that similar subcriticality conditions were necessary for the existence of rings and the Spot B solution in Theorems 4.2 and 4.3, respectively. The difference is that in the SHE the centralizing bifurcation is of Hopf type which leads to the rolling Turing patterns in space. In both cases it is the subcriticality that leads to localized pattern formation. The condition $\mu>\alpha\omega$ provides that solutions in the far-field decay monotonically to zero as $r\to\infty$, leading to the monotonic tails of solutions coming from Theorem 4.4.
@@ 5.2 Localized Patches Beyond Onset
Here they find that the patches appear to undergo snaking behavior, as shown in Figure 36. However, the branch eventually terminates once the patch gets sufficiently large with various isolated branches of larger patches found nearby, as shown in Figure 37. Numerical investigations strongly suggest that the various branches of the localized stripes with $n$ interior rolls should approach vertical asymptotes given by 1D pulses with vanishing Hamiltonian that consist of $n$ rolls, but the reason for this remains a mystery.

### v2

@@ 2.4.3 Slanted Snaking and Finite Domains
We also briefly mention mass conserving systems that also display localized patterns (see for instance [47, 189] where there has been a considerable amount of analysis carried out.
@@ 2.4.4 Oscillons and Breathers
While all the structures we have looked at so far are stationary, time-periodic spatially localized structures have been extensively studied in a wide-range of models. These structures are called either oscillons or breathers, depending on the context they arise in. Oscillons can arise in nonlinear systems from a uniform background state under parametric forcing whose frequency is often close to the frequency of the pattern oscillation in space (the wavenumber of the SPP) [5, 6, 169]. Such localized time-periodic patterns have been observed in Newtonian fluids [11, 105, 249, 269], chemical reactions [202, 245, 243], granular layers [238], and colloidal suspension [157].
@@
In the framework of (2.27), oscillons correspond to localized solutions that are time-independent. The work [58] demonstrated that there are two different types of localized steady-states in (2.27), as depicted in Figure 14. The first is the usual localized solution that limits to $U=0$ as $x\to\pm\infty$. The second limits to a nontrivial background state $U_{\mathrm{unif}}\neq 0$ as $x\to\pm\infty$ and were referred to as reciprocal oscillons, also superoscillons in earlier work [33]. Both of these types of oscillons can have either decay monotonically to $U=0$ as $x\to\pm\infty$, leading to spatial profiles resembling Gaussians/solitons, or damped oscillations into $U=0$, leading to spatial profiles resembling the localized SPPs of the SHE.
@@ 2.4.5 Transition to the Singular Limit
Figure 16: Spatial eigenvalues of the quiescent state for the Belyakov–Devaney transition showing the double-real-eigenvalue transition. We note that this is not a local bifurcation, as no eigenvalues cross the imaginary axis.
@@ 2.4.6 Invasion Fronts Outside the Snaking Region
Figure 17: Outside of the snaking parameter regime localized SPPs are dynamic, growing for $\mu$ to the left (red) of the snaking branches and collapsing in on itself to the right (blue). This growth is non-uniform, alternating between fast and slow expansions, depending on if the pattern is close to a steady-state at a nearby saddle-node.
Throughout this review we primarily focus on steady-state patterns. However, some work has been done to understand dynamic localized patterns. For example, outside of the snaking parameter regime in the SHE one can find that either the SPPs expand into the background state or collapse upon themselves as time progresses, as illustrated in Figure 17. Despite the snaking region of parameter space typically being much smaller than the invasion regions, significantly less is known about these patterns and processes.
The spreading of SPPs into the trivial state is fundamentally different in nature to the spreading of the trivial state. That is, in the case where the trivial state invades the SPP patterns, the speed of the invasion is dictated by the wavelength of the SPPs themselves. However, for the SPP invasion process, since we are in the bistable region, the wavelength of the deposited pattern and the time between stripes being deposited are selected [161, 9]. The SPP invasion fronts do not invade at a constant speed and instead have a “helical” invasion process, coming as a more complicated version of the “bottlenecks” one encounters in parameter regimes near a saddle-node bifurcation [223, Section 4.3]. Precisely, the non-constant traveling speed of the front is understood by looking at the fronts near the snaking region, where one would expect the invasion process to slow down as one passes near a fold bifurcation and then speed up between the folds, as illustrated in Figure 17. There has been a lot of recent progress on the analysis of pattern-forming invasion fronts; see [108] for a review.
@@ 3 Localization in Only One Direction
We begin this section in §3.1 with localized stripe patterns wherein the localized pattern takes the form of a 1D SPP extended trivially in either the direction of $x$, $y$, or a combination thereof. §3.2 reviews the work on patterns with hexagonal patches arranged along the region of localization, while §3.3 provides the spatial dynamics interpretation of these stripe and front patterns. Finally, in §3.4 we review similar work on planar lattices.
@@ 3.3 Spatial Dynamics
Thus, we see that many of the same ingredients that were exploited in [30] (and summarized in §2.3) are present to prove the existence of localized SPPs in 1D. Much like the 1D setting, localized structures of (3.1) correspond to orbits homoclinic to $\mathbf{0}\in\mathcal{X}$, i.e. $\mathbf{u}(x,\cdot)\to\mathbf{0}$ as $x\to\pm\infty$. One needs to be careful here though as (3.6) is ill-posed as an initial-value problem. However, the theory developed in [201, 213] guarantees that stable and unstable manifolds exist for the quiescent state $\mathbf{0}\in\mathcal{X}$, providing the necessary phase-space structure for homoclinic orbits to exist. Beyond this, first recall from §2.3 that spatial dynamics interprets localized solutions as homoclinic orbits wrapping a periodic orbit to produce the localized SPPs. The same interpretation now holds here where the localized pattern comes from the homoclinic orbit of (3.6) wrapping around a periodic in $x$ solution $\gamma(x;y)$. Again, the theory in [213] guarantees the existence of stable and unstable manifolds to $\gamma(x;y)$ in the phase-space $\mathcal{X}$ that can be used to prove the existence of a homoclinic orbit that wraps around $\gamma$ (in $x$), producing the desired localized solution to (3.1).
@@ 4.1 Emergence of Axisymmetric Patterns
Theorem 4.1 (Spot A, [158, Theorem 2]).
@@
Theorem 4.2 (Rings, [158, Theorem 1]).
@@
Notice that these solutions bifurcate from $\mu=0$ in pairs, giving that rings arise via a pitchfork bifurcation from the Turing instability in the SHE. Moreover, the condition $|\nu|>\sqrt{27/38}$ arises from (4.6) and translates to the requirement that the SPPs (Turing patterns) in the SHE bifurcate subcritically. In what follows we refer to solutions along the bifurcating $u_{R}(r)$ branch as Ring A and the solutions along the $-u_{R}(r)$ branch as Ring B.
@@
Theorem 4.3 (Spot B, [177, Theorem 1]).
@@ 4.2 Axisymmetric Snaking Branches
Continuing the axisymmetric patterns of the previous subsection from $\mu\approx 0$ into larger values of $\mu$ reveals a complex bifurcation structure that bears little resemblance to the familiar 1D snakes and ladders profiles [176]. The situation appears to be qualitatively the same for both dimensions $n=2,3$ and Figure 29 presents a numerically obtained bifurcation diagram for visual reference. Initially the Spot A and Spot B branches follow the usual snaking structure wherein fold bifurcations flip back and forth over a compact region of parameter space with the folds appearing to align along two vertical asymptotes. Moreover, like the 1D snaking phenomenon, SPPs are added that extend the region of localization for the axisymmetric pattern, in turn causing the $L^{2}$-norm of the solution to increase, thus resulting in the vertical ascension in the bifurcation diagram.
@@
This vertical growth along the bifurcation diagram is not long lived. Eventually both the Spot A and Spot B branches reach a maximal finite height in the bifurcation diagram where they turn into rings and begin decreasing in norm by subtracting SPPs. Following [176], Spot A connects to the Ring A branch and Spot B the Ring B branch, as defined in the previous subsection. Above both of these terminal snaking branches in the bifurcation diagram lie stacks of isolas - closed curves in the bifurcation diagram. Along each isola the solution profiles change in intricate ways and switch between both ring- and spot-like patterns. In both 2D and 3D there are only finitely many isolas with the 2D patterns extending vertically in the bifurcation diagram until they have approximately 38 SPPs, while in 3D the isolas terminate with fewer SPPs. Finally, above the isolas there is another snaking branch which consists of two intertwined arms that appear to ascend vertically in the bifurcation diagram without bound. This snaking branch is referred to as the upper snaking branch, while the branch containing Spot A or Spot B patterns that emerges from $(U,\mu)=(0,0)$ is termed the lower snaking branch. The upper branch differs significantly from the lower branch in that SPPs are added near the core of the pattern, i.e. near $r=0$, as opposed to at the tail like in 1D. Based on numerical observations, McCalla and Sandstede further conjectured in [176] that the upper branch collapses onto a single parameter value as the number of SPPs in the localized pattern goes to infinity. Bramburger et al. [43] attempted to better understand the bifurcation structure of localized axisymmetric patterns in the two- and three-dimensional SHE from an analytical perspective. As in the previous section, their work focused on the radial spatial dynamical system (4.3). When $n=1$ the nonautonomous terms disappear, reverting back to the 1-dimensional setting (2.10) where all dynamics take place inside the level sets of the Hamiltonian function (2.18). Importantly, when $n=1$ SPPs of the SHE (4.1) correspond to periodic orbits of (4.3) which is only a slice of a ‘stack’ of periodic orbits forming a normally hyperbolic invariant manifold $\mathcal{P}$ that can be parameterized by the value of $\mathcal{H}$. The result is a cylinder of stacked periodic orbits in phase-space, as illustrated in Figure 30. Due to the conserved quantity $\mathcal{H}$, localized SPPs of the Swift–Hohenberg equation in 1D can only be constructed through connections with the periodic orbit in $\mathcal{H}^{-1}(0)$, the zero level set of $\mathcal{H}$, to the rest state $U=0$ for each fixed $(\mu,\nu)$.
@@
Figure 30: Left: The anatomy of a radial SPP solution used in the proofs of [43]. Right: In the phase-space of (4.3) localized radial SPPs wrap around an invariant cylinder in phase-space for a finite time to produce the SPP.
@@
To better understand the system (4.3) (and in turn axisymmetric localized patterns), McCalla and Sandstede proposed introducing a “dimensional perturbation” in the form of $n=1+\varepsilon$ [176]. Although meaningless in the context of the dimension of the underlying space, when $0<\varepsilon\ll 1$ this dimensional perturbation casts (4.3) as a true perturbation from the well-understood spatial system (2.10) in 1D. In particular, [43, Theorem 2.7] shows that snaking branches persist for plateau lengths $L\leq\mathrm{exp}(b/\varepsilon)$, with constant $b>0$. Such an upper bound is the result of determining the maximal number of times a trajectory can be guaranteed to wrap around the perturbed cylinder in complete generality. Surprisingly, numerical continuations of localized radial solutions to (4.3) at fixed values of $\varepsilon$ reveal that the upper bound $\mathrm{exp}(b/\varepsilon)$ closely tracks with the maximum of the lower snaking branches, even well beyond the asymptotic regime $0<\varepsilon\ll 1$ and into $\varepsilon=1,2$ (dimensions $n=2,3$).
@@
There are a number of questions that remain outstanding related to the bifurcation curves of localized radial patterns. Primary among them is that there is still no explanation for the intermediate stack of isolas that splits the lower and upper snaking branches. Similarly, it is not clear why the upper branch forms a connected curve or if it is even the uppermost branch in the bifurcation diagram. That is, it is entirely possible that the observed upper branch terminates much like the lower branch at a finite height in the bifurcation diagram, possibly giving way to more isolas or another snaking branch. Such a finding would not necessarily be surprising as the work [158] originally presented evidence that the lower branch extends indefinitely and it wasn’t until [176] that the termination of the lower branch was observed far up in the bifurcation diagram. Finally, the reason for the difficulty in predicting the specific shape of the snaking diagram with $n=2,3$ is that numerical investigations indicate that the structure is driven by the behavior of solutions away from the invariant manifold of periodic orbits. This means that the curves depend on the global dynamics of the dynamical system (4.3) and not just the local properties near the perturbed cylinder.
@@ 4.3 Time Periodic Patterns
Theorem 4.4 (Radial oscillons, [179, Theorem 1]).
@@
Figure 31: Spatial eigenvalues for the reversible Takens–Bogdanov instability.
This result again makes use of radial spatial dynamics and provides the existence as emergence from a bifurcation. In the setting of (4.10), the curve $\gamma=\sqrt{\mu^{2}+\omega^{2}}$ corresponds to a reversible Takens–Bogdanov bifurcation in the associated radial spatial dynamical systems; see Figure 31. The condition $\mu<\beta\omega$ provides that the bifurcation is subcritical, while we recall that similar subcriticality conditions were necessary for the existence of rings and the Spot B solution in Theorems 4.2 and 4.3, respectively. The difference is that in the SHE the centralizing bifurcation is of Hopf type which leads to the Turing patterns in space. In both cases, it is the subcriticality that leads to localized pattern formation. The condition $\mu>\alpha\omega$ provides that solutions in the far-field decay monotonically to zero as $r\to\infty$, leading to the monotonic tails of solutions coming from Theorem 4.4.
@@ 5.2 Localized Patches Beyond Onset
Here they find that the patches appear to undergo snaking behavior, as shown in Figure 36. However, the branch eventually terminates once the patch gets sufficiently large with various isolated branches of larger patches found nearby, as shown in Figure 37. Numerical investigations strongly suggest that the various branches of the localized stripes with $n$ interior SPPs should approach vertical asymptotes given by 1D pulses with vanishing Hamiltonian that consist of $n$ SPPs, but the reason for this remains a mystery.